Silven Industries, which manufactures and sells a highly successful line of summer lotions and insect repellents, has decided to diversify in order to stabilize sales throughout the year. A natural area for the company to consider is the production of winter lotions and creams to prevent dry and chapped skin.

After considerable research, a winter products line has been developed. However, Silven’s president has decided to introduce only one of the new products for this coming winter. If the product is a success, further expansion in future years will be initiated.

The product selected (called Chap-Off) is a lip balm that will be sold in a lipstick-type tube. The product will be sold to wholesalers in boxes of 24 tubes for $10 per box. Because of excess capacity, no additional fixed manufacturing overhead costs will be incurred to produce the product. However, a $115,500 charge for fixed manufacturing overhead will be absorbed by the product under the company’s absorption costing system.

Using the estimated sales and production of 165,000 boxes of Chap-Off, the Accounting Department has developed the following manufacturing cost per box:

Direct material $ 4.70
Direct labor 3.00
Manufacturing overhead 2.10
Total cost $ 9.80

The costs above relate to making both the lip balm and the tube that contains it. As an alternative to making the tubes for Chap-Off, Silven has approached a supplier to discuss the possibility of buying the tubes. The purchase price of the supplier's empty tubes would be $2.00 per box of 24 tubes. If Silven Industries stops making the tubes and buys them from the outside supplier, its direct labor and variable manufacturing overhead costs per box of Chap-Off would be reduced by 10% and its direct materials costs would be reduced by 30%.

Required:

1. If Silven buys its tubes from the outside supplier, how much of its own Chap-Off manufacturing costs per box will it be able to avoid? (Hint: You need to separate the manufacturing overhead of $2.10 per box that is shown above into its variable and fixed components to derive the correct answer.)

2. What is the financial advantage (disadvantage) per box of Chap-Off if Silven buys its tubes from the outside supplier?

3. What is the financial advantage (disadvantage) in total (not per box) if Silven buys 165,000 boxes of tubes from the outside supplier?

4. Should Silven Industries make or buy the tubes?

5. What is the maximum price that Silven should be willing to pay the outside supplier for a box of 24 tubes?

6. Instead of sales of 165,000 boxes of tubes, revised estimates show a sales volume of 203,000 boxes of tubes. At this higher sales volume, Silven would need to rent extra equipment at a cost of $70,000 per year to make the additional 38,000 boxes of tubes. Assuming that the outside supplier will not accept an order for less than 203,000 boxes of tubes, what is the financial advantage (disadvantage) in total (not per box) if Silven buys 203,000 boxes of tubes from the outside supplier? Given this new information, should Silven Industries make or buy the tubes?

7. Refer to the data in (6) above. Assume that the outside supplier will accept an order of any size for the tubes at a price of $2.00 per box. How many boxes of tubes should Silven make? How many boxes of tubes should it buy from the outside supplier?
Answer #1

Variable manufacturing overhead per box = Total Manufacturing overhead - Fixed manufacturing overhead

= 2.1 - (115500/165000) = 1.4

Total variable cost per box = 4.7 + 3 + 1.4 = 9.1

If tubes purchased from outside supplier then reduced cost will be :-

Cost Reduction in cost Reduced cost
Direct material 4.7 30% 3.29
Direct labor 3 10% 2.7
Variable manufacturing overhead 1.4 10% 1.26
Purchase price of tube 2 0% 2
9.25

1) If Silven buys its tubes from the outside supplier, manufacturing costs = 3.29+2.7+1.26 = 7.25 whereas if it manufactures itself then manufacturing costs = 4.7+3+1.4 = 9.1.(but it will have to pay $2 per box extra for buying tubes).

Therefore it can avoid manufacture cost of $1.85 per box. (If we consider that $2 per box, then it will result in loss of $0.15 per box).

2) Financial disadvantage per box of Chap-Off if Silven buys its tubes from the outside supplier = 9.25 - 9.1 = $0.15

3) Financial disadvantage in total (not per box) if Silven buys 165,000 boxes of tubes from the outside supplier

= (9.25-9.1)*165000 = $24750

4) Silven industries should make the tubes.

5) Maximum price that Silven should be willing to pay the outside supplier for a box of 24 tubes

= Variable cost of producing the tubes - Change in variable cost when buying from outside supplier

= 9.1 - (3.29+2.7+1.26)

=$1.85

6)  

Making Buying
Total relevant cost 1917300 1877750

Making = (203000*9.1)+70000 = 1917300

Buying = 203000*9.25 = 1877750

Silven should buy the tubes as the cost of buying is lower.

7) Alternative 1 :

Units Rate Cost
Cost of making 165000 9.1 1501500
Cost of making 38000 9.1 345800
Rent 70000
203000 1917300

Alternative 2 :

Units Rate Cost
Cost of making 165000 9.1 1501500
Cost of buying 38000 9.25 351500
203000 1853000

Alternative 2 is better as it is less costlier.

Hence we will make 165000 units and buy 38000 units from outside supplier.